OpenStudy (anonymous):

Find the value of m: -9m = 5. m = -9/5 m = 5/9 m = -5/9 m = 9/5

OpenStudy (anonymous):

is it the last answer?

OpenStudy (anonymous):

coz just move it over

OpenStudy (anonymous):

multiply both sides of the original equation by -1/9. What happens?

OpenStudy (anonymous):

it turns +?

OpenStudy (anonymous):

notice that on the left you have (-1) * 9 * m. You want to multiply by something so that you are left with only m on the left. Whatever you multiply by on the left needs to also be done to the right. To preserve the equality.

OpenStudy (anonymous):

so waht will the answer be?

OpenStudy (anonymous):

So first of all you might want to remove the negative on the left. You need to think what you can multiply by to get rid of the negative. Then you need to multiply by something to get rid of the 9. What can you multiply 9 by so that you get 1 (i.e. you "get rid of the 9") ? Then you've gotten rid of the -1 and the 9 on the left. what's left? What's on the right?

OpenStudy (anonymous):

You can't have people just tell you the answer to this. You have to learn how to solve this type of equation. It is quite simple and once you learn you won't need to ask people all the time.

OpenStudy (anonymous):

have you gone through what I suggested? If you do then you'll get the answer...

OpenStudy (anonymous):

this is sooo mind bottling

OpenStudy (anonymous):

It's not. You can pick any number and multiply both sides by that. What number are you going to pick so that m is alone on the left, and on the right is some other number? You need to find a number so that x * (-9) = 1.

OpenStudy (anonymous):

ohhh i get it its 9/5 thx

OpenStudy (anonymous):

\[-9m=5\] \[{-9m \over -9}=-{5\over9}\]

OpenStudy (anonymous):

m is not equal to 9/5.

OpenStudy (anonymous):

yea i know now

OpenStudy (anonymous):

(-1/9) * ( -9 )m = 1 * m = (-1/9) * 5 = -5/9 as monokerous did about. -9 = -9/1. So 1/(-9) * (-9/1) = 1.... remember that and you'll do well.